OpenStudy (anonymous):

The difference between five times a number and twelve times the number is –35.

OpenStudy (solomonzelman):

write it out 5x+35=12x 35=7x x-5

OpenStudy (solomonzelman):

X=5 i meant

OpenStudy (anonymous):

Let x = the number 5x - 12x = -35 -7x = -35 x = 5